OpenStudy (anonymous):

i need help on my math please... QUADRATIC FORMULA i need help on my math please... QUADRATIC FORMULA @Mathematics

OpenStudy (anonymous):

\[x ^{2}-4x+3=0\]

OpenStudy (anonymous):

this acutally factors to (x-1)(x-3) so x = 1, 3

OpenStudy (anonymous):

i need to see the work for quadratic formula please :/

OpenStudy (lilg132):

-1 -3 (x - 1) (x - 3) = 0 x^2 -3x - x + 4 = 0 x^2 - 4x + 4 = 0 x-1 = 0 x - 3 = 0 x = 1 x = 3

OpenStudy (anonymous):

thanks

OpenStudy (lilg132):

dont forget to click good answer :)